# Collective Variable Free Transition Path Sampling with Generative Flow Network

### v1

@@ 2.3 Generative flow networks
Generative flow networks (GFlowNets) are the learning framework for amortized inference, generating objects through a series of decisions with a stochastic policy, or forward policy. [15, 16]. By learning the policy to sample an object proportionally to their rewards, GFlowNets can produce diverse samples [15]. To be specific, they train the forward policy by matching its trajectory flow with an auxiliary backward trajectory flow. Training objectives such as flow matching [15], trajectory balance [17], and detailed balance [16] have been proposed. Recent studies have successfully applied GFlowNets to tasks including Boltzmann generators [33], molecular conformer generation [34], crystal structure design [35], and biological sequence design [36].
@@ 3.2 TPS-GFN: GFlowNets for transition path sampling
We model the bias potential $b_{\theta}(\bm{r})$ or directly bias force $\mathbf{b}_{\theta}(\bm{r})=\nabla_{\bm{r}}b_{\theta}(\bm{r})$ for the forward policy. We use a multi-layered perception (MLP) architecture to parameterize the bias potential or force.
@@ 4.2 Quantitative results
Polyproline Helix and Chignolin. Unlike Alanine Dipeptide, the CVs of Polyproline and Helix, and Chignolin are less understood. Therefore, we exclude metrics THP, ETP, and EFP which require CVs to define the meta-stable state. In Table 2, our method consistently outperforms baselines in both Polyproline Helix and Chignolin, regardless of predicting bias potential or force. Our model samples more realistic transition paths than the baselines with respect to the overall path and final conformation. We note that as the size of the molecule is larger, the EPD and EPCD performances decreases. For the large size of molecules including Chignolin, the bias potential produces better transition paths than bias force since the output of bias potential is always scalar regardless of the size of the molecule and the external force from the bias potential is conservative.
@@ 4.3 Qualitative results
Figure 4: Visualization of a transition path from the meta-stable region PPII to PPI of Polyproline generated by our TPS-GFN. The top row visualizes the 3d view of states along the time horizon. The bottom row visualizes the potential energy along the time horizon. The backbone of the Polyproline is highlighted in green.
@@ Appendix A Generative Flow Networks
Generative flow networks. GFlowNet [15, 16] is a framework to sequentially construct samples $s\in\mathcal{S}$ proportional to a reward function $R(s)$, where $\mathcal{S}$ is a set of data. In other words, GFlowNet aims to learn a target distribution $p_{\ast}(s)$ defined as:
@@
where $Z_{\theta}\in\mathbb{R}$ is a learnable scalar parameter to estimate the normalization constant $Z$. Equation 4 induces $Z_{\theta}P_{F}(s)\approx R(s)$, by marginalizing forward and backward flows over set of trajectories $\mathcal{T}(s)$ ending at the terminal state $s$, i.e., $\int_{\tau\in\mathcal{T}(s)}Z_{\theta}P_{F}(\tau;\theta)=Z_{\theta}P_{F}(s)$ and $\int_{\tau\in\mathcal{T}(s)}R(s)P_{B}(\tau|s;\theta)=R(s)$. In other words, we can sequentially construct a sample $s$ proportional to a reward function $R(s)$ through a series of state transitions by the forward policy after minimizing the flow matching objective.
@@ B.3 Evaluation metrics
$\operatorname{EPD}(\bm{r}_{t},\bm{r}_{\mathcal{B}})=\frac{1}{N^{2}}\lVert D(\bm{r}_{t})-D(\bm{r}_{\mathcal{B}})\rVert^{2}$
@@
$C(r)_{ij}=\begin{cases}0.5Z_{i}^{2.4}&\text{for }i=j\\ Z_{i}Z_{j}/\lVert r_{i}-r_{j}\rVert^{2}&\text{for }i\neq j\end{cases}$

### v2

@@ 2.3 Generative flow networks
Generative flow networks (GFlowNets) are the learning framework for amortized inference, generating objects through a series of decisions with a stochastic policy, or forward policy. [15, 16]. By learning the policy to sample an object proportionally to their rewards, GFlowNets can produce diverse samples [15]. To be specific, they train the forward policy by matching its trajectory flow with an auxiliary backward trajectory flow. Training objectives such as flow matching [15], trajectory balance [17], and detailed balance [16] have been proposed. Recent studies have successfully applied GFlowNets to tasks including molecular conformer generation [33], drug discovery [34], crystal structure design [35], and biological sequence design [36].
@@ 3.2 TPS-GFN: GFlowNets for transition path sampling
We model the bias potential $b_{\theta}(\bm{r})$ or directly bias force $\mathbf{b}_{\theta}(\bm{r})=\nabla_{\bm{r}}b_{\theta}(\bm{r})$ for the generator. We use a multi-layered perception (MLP) architecture to parameterize the bias potential or force.
@@ 4.2 Quantitative results
Polyproline Helix and Chignolin. Unlike Alanine Dipeptide, the CVs of Polyproline Helix, and Chignolin are less understood. Therefore, we exclude metrics THP, ETP, and EFP which require CVs to define the meta-stable state. In Table 2, our method consistently outperforms baselines in both Polyproline Helix and Chignolin, regardless of predicting bias potential or force. Our model samples more realistic transition paths than the baselines with respect to the overall path and final conformation. We note that as the size of the molecule is larger, the EPD and EPCD performances decreases. For the large size of molecules including Chignolin, the bias potential produces better transition paths than bias force since the output of bias potential is always scalar regardless of the size of the molecule and the external force from the bias potential is conservative.
@@ 4.3 Qualitative results
Figure 4: Visualization of a transition path from the meta-stable region PP-II to PP-I of Polyproline generated by our TPS-GFN. The top row visualizes the 3d view of states along the time horizon. The bottom row visualizes the potential energy along the time horizon. The backbone of the Polyproline is highlighted in green.
@@ Appendix A Generative Flow Networks
Generative flow networks. GFlowNet [15, 16] is a framework to sequentially construct samples $s\in\mathcal{S}$ proportional to a reward function $R(s)$, where $\mathcal{S}$ is a set of data. In other words, GFlowNet aims to learn a target distribution $p^{\ast}(s)$ defined as:
@@
[…]
@@ B.3 Evaluation metrics
$\operatorname{EPD}(\bm{r}_{t},\bm{r}_{\mathcal{B}})=\frac{1}{N^{2}}\lVert D(\bm{r}_{t})-D(\bm{r}_{\mathcal{B}})\rVert_{F}^{2}$
@@
$C(r)_{ij}=\begin{cases}0.5Z_{i}^{2.4}&\text{for }i=j\\ Z_{i}Z_{j}/\lVert r_{i}-r_{j}\rVert&\text{for }i\neq j\end{cases}$